OpenStudy (anonymous):

Help Please!! medal! Which of the following best describes the graph of f(x) = x2 - 3x - 10? (4 points) Minimum at (1.5, -12.25) with intercepts at (5, 0) and (-2, 0) Minimum at (-1.5, -12.25) with intercepts at (-5, 0) and (2, 0) Minimum at (-3.5, -2.25) with intercepts at (-5, 0) and (-2, 0) Minimum at (3.5, -2.25) with intercepts at (5, 0) and (2, 0)

OpenStudy (welshfella):

x^2 - 3x - 10 = (x - 1.5)^2 - 12.25

OpenStudy (anonymous):

thank you

OpenStudy (welshfella):

this will give you the coordinates of the minimum value

OpenStudy (welshfella):

you can find the x intercepts by solving x^2 - 3x - 10 = 0

OpenStudy (welshfella):

need any more help?